OpenStudy (anonymous):

Suppose a couple will continue having children until they have at least two children of each sex (two boys and two girls). How many children might they expect to have? For @inkyvoyd @Rohangrr

OpenStudy (anonymous):

Interesting... an expected value problem with infinite outcomes...

OpenStudy (anonymous):

Let \(X\) be the number of children they have. Clearly: \[ \Pr(X<2) = 0 \]So we consider... \(2\leq X\leq \infty\)

OpenStudy (anonymous):

We want to find \[ \Pr (X=n) \]Consider if they have 4 children...

OpenStudy (anonymous):

3 girls and a boy or 3 boys and a girl...

OpenStudy (anonymous):

Now order does matter... so we have: \[ \Pr(X=4) = (0.5)^3(0.5)+(0.5)^3(0.5) = (0.5)^3 \]It seems like our general formula is going to be: \[ \Pr(X=n)=(0.5)^{n-1} = \frac{1}{2^{n-1}} \]

OpenStudy (anonymous):

this would leave our expected value at: \[ \Large E[X] = \sum_{n=2}^{\infty} n\Pr(X=n) =\sum_{n=2}^{\infty} \frac{n}{2^{n-1}} \]